A UNIFORM METER RULE IS PIVOTED AT P IN 70cm mark on the meter rule force 0.1 And 0.4N Are Applied At M The 60Cm Mark And Q The 85Cm Mark. If The Rule Is In Equilibrium .Find The Mass Of The Rule.Take g=10ms-2
1
Expert's answer
2019-02-18T14:32:08-0500

Taking the moments about pivot:

W(0.7-0.5)+0.1(0.7-0.6)=0.4(0.85-0.7)

W=0.25 N.

Thus, the mass of the rule:

m=W/g=0.25/10=0.025 kg=25 g.